# Leakage and Interpretability in Concept-Based Models

### v1

@@ 8 CEMs are not interpretable
CEMs (Espinosa Zarlenga et al., 2022) have gained significant recognition in the domain of interpretable concept-based modelling. However, due to their inherent design these models present fundamental interpretability issues. The primary cause of this is the fact that their reasoning is based on data point-dependent concept vector representations, and not directly on concept activations as in CBMs.
@@
Consider a dataset where the annotated concepts are binary. [^1] The following argument straightforwardly extends to categorical and continuous annotated concepts. A transparent concept-based model trained on such a dataset must generate task predictions solely based on the activation levels of concepts (essentially whether each of them is on or off), which can be supervised during training and later by a human agent. The vector representations $\bm{\hat{c}^{w}}_{i}$ on which CEMs reasoning is based are instead high-dimensional objects capable of encoding an enormous amount of information besides concept activations. Therefore, the necessary condition for CEMs to be deemed interpretable is that the learnt embeddings must be faithful representations of concepts, corresponding to the situation where the generated vectors $\bm{\hat{c}^{w}}_{i}$ essentially contain the same information as the annotated concepts (were these binary, categorical or continuous). [^2] Note that there would be limited advantages in using vector representations if that were the case and they were hence just as informative as concept activations. Besides introducing extra computational overhead during training and inference, the fact that the final head takes as input the concatenated vectors $\bm{\hat{c}^{w}}_{i}$ limits interpretability of the $y=f(c)$ functional dependence, even when interpretable final heads such as a linear layer are used. The final head reasons at the level of the each vector component, which are not interpretable, and statements that can be made are of the form, the $m$-th component of the $i$-th concept vector is more relevant for the task prediction than the $n$-th component of the $j$-th concept vector.
@@
However, in this section we demonstrate that the learnt embeddings in CEMs are not a faithful representation of concepts and that they encode large amounts of leaked information. Intuitively, this is a consequence of no direct supervision being performed on the actual input of the final head, but only on the predicted concept activations $\hat{c}_{i}$ via the concept reconstruction loss. This means that in the forward pass, an arbitrary amount of information is allowed to flow undisturbed from the input $x$ to the task prediction through the concept vectors, as long as the embeddings $\bm{\hat{c}^{+}}_{i}$ and $\bm{\hat{c}^{-}}_{i}$ generated from $x$ are predictive of the single scalar $c_{i}$. Information is free to leak into the learnt vectors, fundamentally obscuring the model reasoning and effectively resulting in a scenario where the model is a black box whose explainability can be only partially recovered using post-hoc methods. This fundamental problem of CEMs is further exacerbated by our findings from Section 7 concerning over-expressive concept representations with respect to the annotated concepts, and in particular how such setups generally favour leakage.
Using our information-theoretic approach, we now show in detail how the learnt concept embeddings are affected by concepts-task leakage for any choice of the hyperparameters. Moreover, we find that contrarily to what happens in CBMs, higher concept supervision actually results in higher interconcept leakage.
Information-theoretic measures to assess CEMs interpretability.
@@
CEMs incorporate several concept representations – the concept probabilities $\hat{c}_{i}$ just as CBMs, as well as the vector encodings $\bm{\hat{c}^{+}}_{i},\bm{\hat{c}^{-}}_{i}$ and $\bm{\hat{c}^{w}}_{i}$. Reasoning happens at the level of the weighted vectors $\bm{\hat{c}^{w}}_{i}$, making them critical objects where leakage could hinder interpretability. As described in Appendix G, entropy and MI estimators are generally affected by biases depending on the dimension of variables. This entails that one cannot compare quantities estimated for concept representations of different dimensions, as required by the definitions of the CTL and ICL scores in equations (6)-(8), which in this case would involve subtracting normalised MIs estimated on the ground-truth concepts $c_{i}$ from those on the vectors $\bm{\hat{c}^{w}}_{i}$. To assess leakage in CEMs, it will be sufficient to analyse the behaviour of the normalised MIs on $\bm{\hat{c}^{w}}_{i}$ across different models without referring to ground-truth normalised MIs, and such a bias will not affect our discussion.
@@
This score captures how informative on average the weighted vectors $\bm{\hat{c}^{w}}_{i}$ are of the task, and it represents a simpler version of the CTL score without the ground-truth concepts-task MI as a reference.
In a similar spirit to the ICL scores, one can further define a metric capturing the amount of information that a concept representation $\bm{\hat{c}^{w}}_{i}$ encodes about the other concepts. Averaged over the $k(k-1)/2$ non-trivial concept pairs, such interconcept normalised MI takes the form,
@@
An auxiliary quantity that in the case of higher-dimensional representations contains non-trivial information is the normalised MI of a concept representation $\bm{\hat{c}^{w}}_{i}$ with the value of the corresponding ground-truth concept $c_{i}$. Its average over the $k$ concepts reads,
@@
providing a measure of how predictive each concept representation is of its own concept.
@@
If the vectors $\bm{\hat{c}^{w}}_{i}$ learnt by CEMs were a faithful representation of concepts, they should be clustered based on the value of the annotated concepts $c_{i}$ only. To test this idea, we train CEMs with a range of $\lambda$ and $p_{int}$, on TabularToy(0.25), dSprites(0) and 3dshapes(0). As we show in Figure 9 (see Figure 25 for similar results at non-vanishing $p_{int}$), this is never the case: although concepts are binary in these experiments, instead of two homogeneous clusters for each concept, we find a rich structure based on the ground-truth value of the task label, for any $\lambda$ and $p_{int}$. This parallels the additional structure appearing in the concept distributions of CBMs affected by leakage, here in the case of a higher dimensional concept representation. This behaviour can be tracked back to the fact that the vectors $\bm{\hat{c}^{w}}_{i}$ are encouraged to be highly predictive of $y$ at any $p_{int}$ and for reasonable values of $\lambda$. Although not perceived there as strong evidence for leakage, a clear embedding structure based on the task label was also reported in Espinosa Zarlenga et al. (2022), Section 5.3.
Information theory offers additional insight into this phenomenon, in particular via the $\widetilde{I}^{(CT)}\left(\bm{\hat{c}^{w}},y\right)$ score defined in (10). As shown in Figure 10, qualitatively its values across models are representative of the degree of refinement in the structure of learnt embeddings displayed in Figures 9 and 25, and furthermore they closely track task accuracy. On the other hand, concept accuracy does not correlate with task accuracy or $\widetilde{I}^{(CT)}$. This not only indicates the presence of concepts-task leakage, but also suggests that CEMs heavily rely on concepts-task leakage for task predictions: regardless of concepts, task accuracy is higher in models where the vectors $\bm{\hat{c}^{w}}_{i}$ encode more information about $y$.
The general dependence of CEMs on leaked information is further substantiated by the fact that task accuracy is insensitive to the number of used concepts, as demonstrated in the experiments reported in Espinosa Zarlenga et al. (2022), Appendix 8. When training CEMs on the CUB dataset with up to 90% of the annotated concepts removed, the task accuracy remains essentially the same as that of models trained on all the concepts. Along the same lines, the rigorous analysis carried out in Espinosa Zarlenga et al. (2022) using the information bottleneck paradigm evidences that in CEMs there is no information bottleneck and that information can freely flow from the input to the task prediction. Initially misunderstood, this result stands as a fundamental proof that the high-dimensional concept representations $\bm{\hat{c}^{w}}_{i}$ generally encode an arbitrary amount of information about the task, meeting the only requirement of being predictive of the real-valued concepts $c_{i}$.
Finally, in Appendix I we report on the subtle effect of alignment leakage, a subtype of concepts-task leakage that may generally appear in models such as CEMs which are exposed to interventions during training. We find evidence of this effect in experiments on the TabularToy($\delta$) dataset. Alignment leakage appears to be stronger in CEMs with higher $\lambda$ and $p_{int}$, supporting its relation with exposure to intervention at training time.
@@
Figure 12: Concept accuracy as well as $\widetilde{I}^{(IC)}$ and $\widetilde{I}^{\text{(self)}}$ scores for CEMs with low and high $\lambda$ and $p_{int}$.
Interconcept leakage increases with stronger concept supervision.
@@
Upon closer inspection of the structure of the learned embeddings $\bm{\hat{c}^{w}}_{i}$, significant clustering is also observed based on the ground-truth values of the other concepts $j\neq i$. This is evidenced in Figure 11, where we display the PCA projections of $\bm{\hat{c}^{w}}_{1}$ in terms of the values of $c_{2}$ and $c_{3}$ (see Figure 26 for similar plots for models with non-vanishing $p_{int}$). As especially manifest in TabularToy(0.25), where the task is binary and there is a total of three concepts, the fine structure present in Figures 9 and 25 which is not explained by the $y$ labels is descriptive of the value of the other concepts.
@@
This can be appreciated more quantitatively using the information-theoretic measures $\widetilde{I}^{(IC)}(\bm{\hat{c}^{w}},c)$ and $\widetilde{I}^{\text{(self)}}(\bm{\hat{c}^{w}},c)$ defined in equations (11)-(12). As indicated in Figure 12 by higher $c_{acc}$ and $\widetilde{I}^{\text{(self)}}$, each vector $\bm{\hat{c}^{w}}_{i}$ becomes more and more predictive of the corresponding $c_{i}$ as one increases $\lambda$ and $p_{int}$. This is the expected and desired behaviour as it corresponds to increasing concept supervision. However, $\widetilde{I}^{(IC)}$ undergoes a very similar growth, entailing that the vectors $\bm{\hat{c}^{w}}_{i}$ contain more and more information about the other concepts. By definition, this represents interconcept leakage. [^3] For this experiment, we deliberately selected datasets where ground-truth interconcept MIs are minimal or vanishing. As a result, there is no interconcept structure for models to learn, nor any justification for concept representations to be highly predictive of the other concepts. In contrast to CBMs, there is no choice of the hyperparameters $\lambda$ and $p_{int}$ that improves concept learning while decreasing interconcept leakage. Although again misinterpreted at the time, concept embeddings were found to be very predictive of the value of the other concepts also in Espinosa Zarlenga et al. (2022), Appendix 9.
@@
Figure 13: CTL and ICL scores evaluated on the predicted concept probabilities $\hat{c}_{i}$ in CEMs with different $\lambda$ and $p_{int}$.
@@
To gain deeper insight into this interconcept leakage in CEMs, we evaluate the CTL and ICL scores on the predicted concept probabilities $\hat{c}_{i}$. Although not indicative of model interpretability – since reasoning occurs at the level of the weighted vectors – it is worth noting that they decrease in a manner similar to CBMs as concept supervision increases. The picture that emerges is thus that as one raises $\lambda$ and $p_{int}$, the leaked information vanishes from the concept probabilities, and is effectively transferred into the vector representations as interconcept leakage. Vector representations in CEMs hence provide a mechanism for interconcept leakage to persist even at high concept supervision.
@@ 9 Lessons for model design
Possible solutions to concept set incompleteness have been studied, including adding unsupervised concepts or channels, meant to account for the missing information (see Mahinpei et al., 2021; Havasi et al., 2022; Sawada and Nakamura, 2022; Yuksekgonul et al., 2023; Sheth and Kahou, 2023; Ismail et al., 2024, 2025). However, when adopting this approach, one has to ensure the model does not over-rely on such unsupervised channels, effectively bypassing the concept bottleneck when making a task prediction. Especially for high-risk scenarios, a general framework to interpret unknown concepts and safely include them in the concept refinement process is still missing.
It is also worth noting that an incomplete concept set and a misspecified classifier head are in principle degenerate issues, both leading to a drop in task performance and similarly higher leakage scores. Quantifying the completeness of a concept set for a given task is by itself an information-theoretic problem that has been extensively discussed in the context of feature selection, and measures have been presented e.g. in Yeh et al. (2020); Havasi et al. (2022), which may provide indications to break such a degeneracy.
@@
For binary annotated concepts, two common setups are soft and hard CBMs. To reduce leakage and achieve accurate concept learning, soft CBMs should be trained at relatively high $\lambda\in(\lambda_{min},\lambda_{max})$, according to the results of Section 7. The $CTL$ and $ICL$ scores enable tuning $\lambda$ to values that minimise leakage. Throughout our experiments they also evidenced that leakage is generally non-vanishing in soft CBMs. This may be acceptable depending on the use case, particularly since at high $\lambda$ the model’s reliance on leaked information may be secondary to concept activations. Evaluating task performance on a number of soft and hard models, we also observe that soft CBMs with high $\lambda$ are typically superior to hard CBMs in terms of task performance only in the cases where the classifier head is misspecified, while being essentially comparable in the other cases (Table 4). Note that when designing a concept-based model, it can be beneficial to also train an end-to-end black-box model without concept supervision, in order to establish an upper bound on the overall task performance and to assess the intrinsic predictivity of the input data for the task.
The results of this work support the information-theoretic framework that we propose as a novel and fundamental paradigm for assessing the interpretability of concept-based models. The leakage measures we introduce are robust and sensitive, and we advocate for their adoption as standard tools for ensuring interpretability when designing concept-based models.
@@
We thank Aya Abdelsalam Ismail, Anthony Baptista and Gregory Verghese for insightful discussions on related topics. All authors except CH are supported by the Turing-Roche strategic partnership. CRSB is supported by the CRUK City of London Centre Award [CTRQQR-2021\100004]. TC is supported by a principal research fellowship from University College London.

### v2

@@ 8 CEMs are not interpretable
CEMs (Espinosa Zarlenga et al., 2022) have gained significant recognition as models that are potentially able to overcome the accuracy-interpretability trade-off. However, due to their inherent design these models present fundamental interpretability issues. The primary cause of this is the fact that their reasoning is based on data point-dependent concept vector representations, and not directly on concept activations as in CBMs.
@@
To see this, consider a dataset in which the annotated concepts are binary (the following argument straightforwardly extends to categorical and continuous annotated concepts). A transparent concept-based model trained on such a dataset must generate task predictions solely based on the activation levels of concepts (essentially whether each of them is on or off), which can be supervised during training and later by a human overseer. However, the vector representations $\bm{\hat{c}^{w}}_{i}$ on which CEMs reasoning is based are high-dimensional objects capable of encoding (a potentially enormous amount of) information besides concept activations. Therefore, a necessary condition for CEMs to be interpretable is that the learnt embeddings faithfully represent the concepts; or, more precisely, the generated vectors $\bm{\hat{c}^{w}}_{i}$ must contain the same information as the annotated concepts.
@@
Here we will show that the learnt embeddings in CEMs are not a faithful representation of concepts and, when properly quantified, that they cannot help but encode large amounts of leaked information. Intuitively, this is a consequence of the fact that no direct supervision is performed on the input to the final head, but only on the predicted concept activations $\hat{c}_{i}$ via the concept reconstruction loss. This means that in the forward pass, an arbitrary amount of information is allowed to flow undisturbed from the input $x$ to the task prediction through the concept vectors, as long as the embeddings $\bm{\hat{c}^{+}}_{i}$ and $\bm{\hat{c}^{-}}_{i}$ generated from $x$ are predictive of the single scalar $c_{i}$. In essence, information is free to leak into the learnt vectors, fundamentally obscuring the model reasoning and effectively producing a black box whose behaviour can only partially be explained using post-hoc methods. This fundamental problem of CEMs is further exacerbated by our findings from Section 7 concerning over-expressive concept representations with respect to the annotated concepts, and in particular how such setups generally favour leakage.
Not only does such leakage occur for any choice of the hyperparameters, but we also find that for CEMs (in contrast to CBMs) higher concept supervision typically results in higher interconcept leakage.
Information-theoretic measures of CEM interpretability.
@@
CEMs incorporate several concept representations – the concept probabilities $\hat{c}_{i}$ just as CBMs, as well as the vector encodings $\bm{\hat{c}^{+}}_{i},\bm{\hat{c}^{-}}_{i}$ and $\bm{\hat{c}^{w}}_{i}$. Reasoning happens at the level of the weighted vectors $\bm{\hat{c}^{w}}_{i}$, making them critical objects where leakage could hinder interpretability. Entropy and MI estimators are generally affected by biases depending on the dimension of variables, as a consequence of the curse of dimensionality (Lord et al., 2018; Carrara and Ernst, 2019; Czyż et al., 2023; Gowri et al., 2024). Thus, one cannot compare quantities estimated for concept representations of different dimensions as required by the definitions of the CTL and ICL scores in equations (6)-(8), since this would involve subtracting normalised MIs estimated on the ground-truth concepts $c_{i}$ from those on the vectors $\bm{\hat{c}^{w}}_{i}$. To avoid this bias we will therefore analyse the behaviour of the normalised MIs on $\bm{\hat{c}^{w}}_{i}$ across different models without referring to ground-truth normalised MIs.
@@
This score captures how informative the weighted vectors $\bm{\hat{c}^{w}}_{i}$ are, on average, of the task. This score can be considered as a simpler version of the CTL score we introduced for CBMs, that does not use the ground-truth concepts-task MI as a reference.
In a similar spirit to the ICL scores, we can also define a simplified metric capturing the amount of information that a concept representation $\bm{\hat{c}^{w}}_{i}$ encodes about the other concepts. Averaging over the $k(k-1)/2$ non-trivial concept pairs, we get:
@@
An associated quantity that is relevant whenever higher-dimensional representations contain non-trivial information is the normalised MI of a concept representation $\bm{\hat{c}^{w}}_{i}$ relative to its ground-truth concept $c_{i}$. Its average over the $k$ concepts is:
@@
which provides a measure of how predictive each concept representation is of its own concept, averaged over concepts.
@@
If the vectors $\bm{\hat{c}^{w}}_{i}$ learnt by CEMs were a faithful representation of concepts, then they should cluster based on the value of the annotated concepts $c_{i}$ only. To test this idea, we trained CEMs with a range of $\lambda$ and $p_{int}$, on TabularToy(0.25), dSprites(0) and 3dshapes(0) data. As we show in Figure 9 (see Figure 21 for similar results at non-vanishing $p_{int}$), this is never the case: although concepts are binary in these experiments, instead of two homogeneous clusters for each concept, we find a rich structure based on the ground-truth value of the task label, for any $\lambda$ and $p_{int}$. This result parallels the additional structure that appears in the concept distributions of CBMs affected by leakage (in this case, accounting for higher dimensional concept representations). This behaviour can be explained by the fact that the vectors $\bm{\hat{c}^{w}}_{i}$ are optimised to be predictive of $y$ at any $p_{int}$, for reasonable values of $\lambda$. Although not interpreted as evidence of leakage, a clear embedding structure based on the task label was also reported in Espinosa Zarlenga et al. (2022).
Information theory offers additional insight into this phenomenon, in particular via the $\widetilde{I}^{(CT)}\left(\bm{\hat{c}^{w}},y\right)$ score defined in (10). As shown in Figure 10 this score is representative of the degree of structure in the learnt embeddings given in Figures 9 and 21, and is closely associated with task accuracy. On the other hand, concept accuracy does not correlate with task accuracy or $\widetilde{I}^{(CT)}$. This not only indicates the presence of concepts-task leakage, but also suggests that CEMs rely heavily on concepts-task leakage for their success: regardless of concepts, task accuracy is higher in models where the vectors $\bm{\hat{c}^{w}}_{i}$ encode more information about $y$.
Building on these observations, in Appendix H we show that alignment leakage, a subtype of concepts-task leakage, can also appear in models such as CEMs which are exposed to interventions during training. We find evidence of this effect in experiments on the TabularToy($\delta$) dataset. Alignment leakage appears to be stronger in CEMs with higher $\lambda$ and $p_{int}$, supporting its relation with exposure to intervention at training time.
@@
Figure 12: Concept accuracy, $\widetilde{I}^{(IC)}$ and $\widetilde{I}^{\text{(self)}}$ scores for CEMs with low and high $\lambda$ and $p_{int}$.
Interconcept leakage increases with concept supervision.
@@
On closer inspection of the structure of the learned embeddings $\bm{\hat{c}^{w}}_{i}$, strong clustering is also observed based on the ground-truth values of the other concepts $j\neq i$. For example, Figure 11 shows PCA projections of $\bm{\hat{c}^{w}}_{1}$ in terms of the values of $c_{2}$ and $c_{3}$ (see Figure 22 for similar plots for models with non-vanishing $p_{int}$). This phenomenon is particularly apparent in the TabularToy(0.25) data, where the task is binary and there are three concepts: in this case, the fine structure present in Figures 9 and 21 which is not explained by the $y$ labels is clearly associated with the value of the other concepts.
@@
This effect can be quantified using the information-theoretic measures $\widetilde{I}^{(IC)}(\bm{\hat{c}^{w}},c)$ and $\widetilde{I}^{\text{(self)}}(\bm{\hat{c}^{w}},c)$ defined in equations (11)-(12). Each vector $\bm{\hat{c}^{w}}_{i}$ becomes more and more predictive of the corresponding $c_{i}$ as one increases $\lambda$ and $p_{int}$, as indicated in Figure 12 by higher $c_{acc}$ and $\widetilde{I}^{\text{(self)}}$. This is the expected and desired behaviour as it corresponds to increasing concept supervision. However, $\widetilde{I}^{(IC)}$ undergoes a very similar growth, indicating that the vectors $\bm{\hat{c}^{w}}_{i}$ contain increasing information about the other concepts, which, by definition, is interconcept leakage. (For this experiment, we deliberately selected datasets where ground-truth interconcept MIs are minimal or vanishing. As a result, there is no interconcept structure for models to learn, nor any justification for concept representations to be highly predictive of the other concepts.) Thus (and in contrast to CBMs) there is no choice of the hyperparameters $\lambda$ and $p_{int}$ that improves concept learning while decreasing interconcept leakage.
@@
Figure 13: CTL and ICL scores evaluated on the predicted concept probabilities $\hat{c}_{i}$ in CEMs with different values of $\lambda$ and $p_{int}$.
@@
To gain deeper insight into this interconcept leakage in CEMs, we evaluated the CTL and ICL scores on the predicted concept probabilities $\hat{c}_{i}$. Although not indicative of model interpretability – since reasoning occurs at the level of the weighted vectors for CEMs – it is worth noting that they decrease in a manner similar to CBMs as concept supervision increases. The picture that emerges is thus that as one raises $\lambda$ and $p_{int}$, the leaked information vanishes from the concept probabilities, and is effectively transferred into the vector representations as interconcept leakage. The vector representations that are used in CEMs hence provide a mechanism for interconcept leakage to persist even at high concept supervision.
@@ 9 Lessons for model design
Possible solutions to concept set incompleteness have been studied, including adding unsupervised concepts or channels, meant to account for the missing information (see Mahinpei et al., 2021; Havasi et al., 2022; Sawada and Nakamura, 2022; Yuksekgonul et al., 2023; Sheth and Kahou, 2023; Ismail et al., 2024, 2025). However, when adopting this approach, one has to ensure the model does not over-rely on such unsupervised channels, effectively bypassing the concept bottleneck when making a task prediction, and becoming an effective black-box. Especially for high-risk scenarios, a general framework to interpret unknown concepts and safely include them in the concept refinement process is still missing.
It is also worth noting that an incomplete concept set and a misspecified classifier head are in principle degenerate issues, both leading to a drop in task performance and similarly higher leakage scores. Quantifying the completeness of a concept set for a given task is by itself an information-theoretic problem that has been extensively discussed in the context of feature selection, and some measures have been presented, for example in Yeh et al. (2020); Havasi et al. (2022), which may provide indications of ways to avoid such degeneracy.
@@
For binary annotated concepts, two common setups are soft and hard CBMs. To reduce leakage and achieve accurate concept learning, soft CBMs should be trained at relatively high $\lambda\in(\lambda_{min},\lambda_{max})$, according to the results of Section 7. The $CTL$ and $ICL$ scores enable tuning $\lambda$ to values that minimise leakage. Throughout our experiments they also indicate that leakage is often unavoidable in soft CBMs. This may be acceptable depending on the use case, particularly since at high $\lambda$ the model’s reliance on leaked information may be secondary to concept activations. By evaluating task performance on a number of soft and hard models, we also observed that soft CBMs with high $\lambda$ are typically superior to hard CBMs in terms of task performance only if the classifier head is misspecified, while being essentially comparable in the all other cases (Table 4). Note that when designing a concept-based model, it can also be beneficial to train an end-to-end black-box model without concept supervision, in order to establish an upper bound on the overall task performance, and assess the intrinsic suitability of the input data for the task.
In summary we have proposed an information-theoretic framework to assessing the interpretability of concept-based models. The leakage measures we introduce are robust and sensitive, and we propose their use as standard tools for ensuring interpretability when designing concept-based models.
@@
We thank Aya Abdelsalam Ismail, Anthony Baptista and Gregory Verghese for insightful discussions on related topics. All authors except CH are supported by the Turing-Roche partnership. CRSB is supported by the CRUK City of London Centre Award [CTRQQR-2021\100004]. TC is supported by a principal research fellowship from University College London. CH is an employee of Roche.